OpenStudy (anonymous):

b-2/2=10/6 7/6=r-3/5

OpenStudy (anonymous):

@Outh-1

OpenStudy (anonymous):

\[b-2/2=10-6\]

OpenStudy (anonymous):

Is that what you are asking for?

OpenStudy (anonymous):

solving proportions

OpenStudy (anonymous):

2/2 simplifies to 1, and 10/6 simplifies to 5/3. That gives you b-1=5/3. You then find the least common multiple of 1 and 5/3, and that is three. You set the denominators both to 3, so b-3/3=5/3, and then you solve it from there.

OpenStudy (anonymous):

Same with the other problem, find the least common multiple, then solve for r.